OpenStudy (mrhoola):

quick verification :

OpenStudy (mrhoola):

\[e ^{-j 25 (rads)}\] i) 25 rads = 25* 180/pi = 1432.394 deg ii) 360 * 3 = 1080 deg iii) 1432.394 - 1080 = 356.394 deg iv) 356.394 deg = -3.606 deg ? There fore:\[e^{j25 rad} = e ^{j -3.606 \deg} \] Is that correct ?

OpenStudy (mrhoola):

\[e ^{j -3.606} = 1 \angle -3.606 \deg\]

OpenStudy (math&ing001):

What are you supposed to prove ?

OpenStudy (irishboy123):

i'd divide 25 by \(\pi\) as many times as necessary to get it under \(2 \pi\) ...so that you can find a sensible complex number

OpenStudy (mrhoola):

analyzing .....

OpenStudy (mrhoola):

@IrishBoy123 I was able to see that 25 rads goes into 7*pi rads 1.1369 times

OpenStudy (mrhoola):

25 * 0.13682 = 3.4202 rads

OpenStudy (irishboy123):

not actually sure what you are trying to show but here are some more data points! a https://www.wolframalpha.com/input/?i=25++rads https://www.wolframalpha.com/input/?i=.1327+*+pi